OpenStudy (anonymous):

please help me, integral of 4 sin at dt is...

OpenStudy (kainui):

Sure, show me your best guess and we'll go from there so you can know exactly what you need to change to be right! =D

OpenStudy (anonymous):

\[\int\limits_{}^{} 4 \ \sin \frac{ n \pi }{ 4} t \]

OpenStudy (anonymous):

how about \[\frac{ - 4 \cos \frac{ n \pi }{ 4 } t }{ \frac{ n \pi }{ 4 } }\] @Kainui ??????

OpenStudy (kainui):

Yep, that's right. =D

OpenStudy (anonymous):

how about this @Kainui \[\int\limits_{}^{} (-\frac { 2 }{ 6 }t) \sin \frac{ n \pi }{ 4 } t = \frac{ \frac{ 2 }{ 6}\cos \frac{ n \pi }{ 4 } }{ \frac{ n \pi }{ 4 }} - \frac{ \frac{ 2 }{ 6 } \cos \frac{ n \pi }{ 4 }t }{({\frac{ n \pi }{ 4}})^2 }\]

OpenStudy (kainui):

Here, since all you really need is someone to confirm if it's right, this might be better than waiting for me or someone else. Then if things don't match up, you can ask for help, but it even offers to show the steps. http://www.wolframalpha.com/input/?i=-2t%2F6sin%28n*pi%2F4t%29dt

OpenStudy (anonymous):

oh, many thankz @Kainui :D

OpenStudy (kainui):

Yeah good luck in your calculus endeavors!

OpenStudy (anonymous):

sure! :D